# Bridging Synthetic and Real Routing Problems via LLM-Guided Instance Generation and Progressive Adaptation

### v1

@@ 1 Introduction
Being a longstanding NP-hard challenge, Vehicle Routing Problems (VRPs) represent a classic family of combinatorial optimization problems (COPs) (BENGIO2021405; cappart2023combinatorial) widely encountered in the field of logistics (CATTARUZZA201751) and public transportation (konstantakopoulos2022vehicle). In recent years, Neural Combinatorial Optimization (NCO) methods have demonstrated remarkable success in tackling various VRPs through deep reinforcement learning and attention-based models, achieving state-of-the-art performance on synthetic VRP instances (bello2017neuralcombinatorialoptimizationreinforcement; kool2018attention; li2021learning; drakulic2023bq; ma2023learning). Despite this progress, existing NCO models frequently exhibit limited generalization when transitioning from synthetic, uniformly distributed training data to real-world problem instances, such as those found in benchmarks like TSPLib (reinelt1991tsplib) and CVRPLib (uchoa2017new). This distributional shift significantly restricts their practical applicability (joshi2006learning; bi2022learning; DBLP:conf/ijcai/GaoS00024).
@@
On the other hand, LLM solvers exhibit strong generalization capabilities when confronted with new COPs and distributions. Recent works leverage LLMs to dynamically generate heuristics or evolve hyper-heuristics (duflo2019gp; drake2020recent), thus facilitating efficient Evolutionary Search (ES) of heuristic spaces without human bias (yang2024large). Noteworthy endeavors include FunSearch (romera2024mathematical), EoH (liu2024evolution), and ReEvo (ye2024reevo). However, existing LLM-based approaches often encounter challenges when handling tasks that require extensive contextual information, typically demonstrating reduced coherence and accuracy in the processing of extended textual descriptions (yang2024large; xu2025can). Consequently, these approaches exhibit significant limitations when addressing medium- or large-size instances (e.g., instances involving more than 50 nodes), thus compromising their ability to effectively generalize to real-world VRP instances.
@@ 2 Related Works
Those models learn policies to construct solutions in a step-by-step or one-shot manner: Pointer Network (Ptr-Net) proposed by (NIPS2015_29921001) employed a transformer-based encoder-decoder network that generates solutions to VRPs in a sequential way. Follow-up works utilize reinforcement learning (RL) to improve Ptr-Net’s gradient update to generate better approximate solutions on TSP (bello2017neuralcombinatorialoptimizationreinforcement) and CVRP (NEURIPS2018_9fb4651c). More recently, more neural VRP models have evolved with their architectures built upon attention mechanisms, starting with the attention-based model (AM) (kool2018attention), which is versatile on a wide range of COPs. Policy optimization with multiple optima (POMO) (NEURIPS2020_f231f210), has further promoted the capability of the attention-based model on TSP and CVRP. Other constructive frameworks developed from AM and POMO illustrate their broader scalability (NEURIPS2021_29539ed9; choo2022simulation; NEURIPS2022_0cddb777; chen2023neural; chalumeau2023combinatorial; hottung2024polynet; ijcai2024p769).
Recent studies have investigated the severe generalization ability decrement in unseen problem sizes or distributions (joshi2006learning; 10188470). In the line of size generalization, there are many attempts to generalize solvers from small instances to larger ones (lisicki2020evaluating; kim2022scaleconditioned; bdeir2022attention; hou2023generalize; pmlr-v202-son23a). Unlike approaches focused on scaling, considerable efforts have been devoted to addressing cross-distribution generalization, including works that augment training instances on multi-distribution and multitasks (goh2025shield), and works which synergetically train a backbone model with multi-distribution instances towards a generalizable solver on out-of-distribution tasks (pmlr-v202-zhou23o; NEURIPS2024_dc709714). Recently, training LLM for end-to-end CO has been attempted, moving beyond prompting paradigms (jianglarge).
@@
Existing approaches to improving generalization in combinatorial optimization (CO) take advantage of the linguistic comprehension and domain-specific competence of LLMs via prompt engineering to tackle COPs. These methods can be mainly divided into two mainstreams: 1) LLMs are solution generators. Early explorations in this line of research involve prompting LLMs to solve graph-based COPs (NEURIPS2023_622afc4e), iteratively refining current heuristics (NEURIPS2024_eb9120be), progressively improving solutions by targeting better objectives (yang2024large; 10611913; make6030093), and instilling LLMs with visual modalities of problems and solutions (elhenawy2024eyeballing; DBLP:journals/corr/abs-2403-01757). 2) LLMs are heuristic designers. Explainable, task-oriented heuristics in the form of codes are iteratively improved throughout automatic heuristic generation and evaluation of Evolution Algorithms (EA), the productivity of which outstrips that of human experts (romera2024mathematical; liu2024evolution; ye2024reevo; tran2025large; jiang2025droc). Among them, ReEvo (ye2024reevo) achieves strong adaptability in many COPs under both white-box and black-box prompt settings. In this paper, we focus mainly on the utilization of LLMs as code designers. Distinct from prior constructive neural solvers and LLM-as-optimizer approaches, our method leverages LLMs to evolve data generators, enabling a novel, generator-based adaptation pipeline that outperforms direct fine-tuning on real data.
@@ 3 Methodology
Representation of Generators. Each generator $h$ is described by three components: 1) a function description defines the format of input and describes valid output, 2) a code implementing generation, and 3) a fitness score. The fitness score, defined as a function $f$ of $h$ ( $f:h\rightarrow\mathbb{R}$), serves as the performance indicator of $h$ on the validation set. An example of an evolved generator that outputs TSP problems is given in Fig.7 in section E of the extended version.
Proxy Evaluation. Following recent work (DBLP:journals/corr/abs-2403-00827), we use an approach that leverages external proxy metrics such as gaps or objective values as the divergence measure $\mathcal{D}$ in Eq.1 to evaluate the quality of LLM-generated outputs, in lieu of direct human or reference-based evaluation. During the LLM-driven evolution, the proxy evaluation is instantiated by measuring the average gap on the validation set, which serves as the proxy metric. The gap is the gap between the objective value found and the best-known optimum. Each generator only fine-tunes the model for a small number of epochs (e.g. 10 epochs for LEHD), and the best average gap on the validation set of each generator during training is taken as the fitness score. Generators with lower fitness scores are retained according to a fixed population size, guiding generator selection without full fine-tuning. This low-fidelity evaluation strategy substantially accelerates the evolution search while presenting reliable indications of the proximity of the generated distributions to the real ones, as is the standard in the neural architecture search and hyperparameter optimization literature.
@@
Evolution Process. In order to balance the effectiveness of evolutionary search and the efficiency of fitness evaluation, we incorporate targeted modifications into ReEvo (ye2024reevo) framework. Ours introduces a ranking-based selection mechanism that increases the survival rate of elite generators, thereby accelerating the search process and reducing token costs. In addition, by postponing the selection stage until after mutation, we balance exploration and exploitation, allowing more newly initialized generators to participate in crossover and mutation before selection. Table 5 in Section 5 shows the elevated performance of the evolution framework equipped with these modifications. We perform the workflow in Fig.2 to search for the best generator by repeating the evolution part for N iterations. The whole evolution procedure can be summarized as the following steps:
@@
Phase One. In the first phase of progressive fine-tuning, the best-performing generator of all types evolved collaboratively generate large-scale synthetic data of the same problem size. By training pre-trained models to the disparate, multi-size node patterns in these diverse, TSPLib-style instances, neural solvers are better equipped to handle challenging patterns in real-world scenarios through preliminary exposure to similar structures. Compared to the proxy evaluation conducted during LLM-driven generator evolution, phase one involves a greater number of training epochs. Additionally, we continuously assess model performance throughout phase one using the 48 validation problems.
@@ 4 Experimental Results
In this section, we empirically validate the effectiveness of EvoReal on two typical VRP benchmarks, TSPLib95 (reinelt1991tsplib) and CVRPLib setX (uchoa2017new), across a wide range of problem sizes.
@@
Baselines. We employ four classic solvers as non-neural baselines: Concorde (applegate2006concorde), ORTools, HGS (VIDAL2022105643) and LKH-3 (helsgaun2017extension). For deep neural solvers, POMO (NEURIPS2020_f231f210), LEHD (luo2023neural), SGBS (choo2022simulation), BQ (drakulic2023bq), ELG (DBLP:conf/ijcai/GaoS00024), CNF (NEURIPS2024_dc709714) and DIFUSCO (sun2023difusco). Importantly, POMO is trained using reinforcement learning (RL) with policy gradient methods, whereas LEHD employs a supervised learning (SL) scheme that leverages optimal objective values. These models sample solely uniform distributions for training.
@@
Training Setups. Hyperparameters have two sets: one for proxy evaluation and the other for progressive fine-tuning. We choose two representative attention-based neural models for our study, namely POMO and LEHD. For both sets, we leverage the evolved generators to produce TSP or CVRP of size 100 for fine-tuning. In phase two, instances from VRPLib are expanded by a pre-set batch size as a single batch, and they are expanded only once in one epoch. For LEHD, Concorde solver and HGS are adopted to obtain the optimum of TSP and CVRP, respectively. For generator evolution, we use OpenAI model o3 to construct new generators. All evolution pipelines can be conducted on a single NVIDIA RTX 3090 Ti GPU with 24GB memory, whereas the progressive fine-tuning is performed on one NVIDIA RTX A5000 GPU with 40GB memory. The complete hyperparameters are given in section D of the extended version.
Metric and Inference. We compare the results with different metrics with respect to different problem size intervals, including average objective values, average gaps (to the optimal average tour length), and total inference time. In order to monitor the performance of models during the proxy evaluation and progressive fine-tuning, we evaluate the models’ performance at fixed epochs. At each validation point, we calculate the total average gap on the validation set as the performance metric with 8-fold augmentation for POMO or greedy mode for LEHD. The inference time of classic non-neural solvers (CPU) and learning-based models (GPU) are not directly comparable. For POMO, LEHD, SGBS, CNF(3), ELG, DIFUSCO (Ts=50) we report the performance of their pre-trained models with their default settings. For BQ, we reproduced their model with their original setups. For the classic non-neural solvers for TSP, we set a maximum time limit according to problem size $n$ ($T_{max}=1.8*n$ seconds for LKH-3 and ORTools, and $T_{max}=10$ minutes for Concorde). As for non-neural solvers for CVRP, we cite the results from Table 2 of HGS paper (VIDAL2022105643). For POMO, we provide the results with and without 8 times augmentation (×8 AUG and non aug). With respect to LEHD, the inference results under both greedy mode and RRC (with 50 iterations) are documented.
@@ 7 Acknowledgments
(a) Histogram of the FFT energy of S1 problems and non-S1 problems in the validation set. The black dotted line represents FFT=35.
(b) Histogram of the NN-ratio of S2 type instances and S3 type instances. The black dotted line represents NN ratio=0.5.
Figure 4: Left: threshold for FFT energy of S1 type and non-S1 type. Right: threshold for NN-ratio of S2 type and S3 type. The thresholds of the divisions are both marked with black dotted line.
(a) Box-plot for FFT energy of S1, S2 and S3 validation sets.
(b) Box-plot for NN-ratio of S1, S2 and S3 validation problems .
Figure 5: Box-plots of FFT energy and NN-ratio of all validation instances after instance segmentation. The corresponding statistics of 5000 TSP 100 instances sampled from uniform distribution is supplemented for comparison.
@@ Appendix A Pre-processing
We follow pre-trained models such as POMO (NEURIPS2020_f231f210) and LEHD (luo2023neural), which randomly sampled uniformly distributed training instances within a unit square. To maintain consistency in data distribution and scale, we also perform the normalization procedure on the node coordinates of TSPLib (reinelt1991tsplib) and CVRPLib (uchoa2017new), rescaling them to the unit square before incorporating these datasets into our training pipeline. The Algorithm 1 is applied to the coordinates of all nodes. We also tried other common techniques like min-max normalization, standardization, and without-normalization for fine-tuning, but they either fail to converge or degrade the performance of the pre-trained model. We empirically find that rescaling TSPLib data to match the scale of synthetically generated uniform data, while preserving the original distributional structure, results in a more favorable convergence landscape during fine-tuning with TSPLib instances. For CVRPLib, we also employ the same approach to normalize customer nodes while re-scale the demands with capacities. For LEHD, we normalize the nodes and other constraints of the TSPLib and CVRPLib to the same scale as their training data.
@@
Several prior studies have used statistics to analyze the structural diversity and heterogeneity of TSP instances (herraiz2022equivalent) or predict optimal tours (kou2022optimal). Motivated by these findings, our approach adopts a structure-driven modularization strategy: we categorize TSPLib instances according to statistical features and design specialized generators tailored to each identified category. Specifically, we partition TSPLib using two complementary statistical measures: Fast Fourier Transform (FFT) energy and Nearest-Neighbor ratio (NN-ratio, also known as the Coefficient of Variation, CV). The formal definitions of these two statistics are as follows:
  • FFT Energy: A quantitative statistic that measures the global periodicity or repetitiveness of a spatial point pattern by evaluating the strength of frequency components on its density map. A high FFT energy indicates pronounced global regularity, periodicity, or repeated structures, while a low FFT energy suggests more disordered, chaotic structures. Inspired by classical spectral analysis in signal and spatial statistics (briggs1995dft; 10107467), FFT energy is widely used to characterize spatial repetition and frequency content in spatial data.
@@
    where $d_{\mathrm{NN}}$ denotes all the nearest neighbor distances in the spatial point set (clark1954distance; illian2008statistical). A low NN-ratio indicates that most points are regularly spaced, as in grid-like or weak-clustered layouts; and a high NN-ratio indicates the presence of clusters or highly varied point spacing. It is a standard, parameter-free measure in spatial statistics to distinguish between regular, random, and clustered point patterns (with self-exclusion, that is, $k=2$ for KNN).
@@
We have also explored other unsupervised machine-learning methods for instance segmentation, such as spectral clustering (von2007tutorial), we found that the classes of most instances (85%) remain the same as those of rule-based segmentation while resulting in trivial performance variation ($5\%$) of the best evolved generators found. This consistency further substantiates the interpretability and robustness of our rule-based classification scheme, indicating that the identified structural discrepancies among different TSPLib distributions are intrinsic rather than heuristic artifacts.

### v2

@@ 1 Introduction
Being a longstanding NP-hard challenge, Vehicle Routing Problems (VRPs) represent a classic family of combinatorial optimization problems (COPs) (Bengio et al. 2021; Cappart et al. 2023) widely encountered in the field of logistics (Cattaruzza et al. 2017) and public transportation (Konstantakopoulos et al. 2022). In recent years, Neural Combinatorial Optimization (NCO) methods have demonstrated remarkable success in tackling various VRPs through deep reinforcement learning and attention-based models, achieving state-of-the-art performance on synthetic VRP instances (Bello et al. 2017; Kool et al. 2019; Li et al. 2021; Drakulic et al. 2023; Ma et al. 2023). Despite this progress, existing NCO models frequently exhibit limited generalization when transitioning from synthetic, uniformly distributed training data to real-world problem instances, such as those found in benchmarks like TSPLib (Reinelt 1991) and CVRPLib (Uchoa et al. 2017). This distributional shift significantly restricts their practical applicability (Joshi et al. 2006; Bi et al. 2022; Gao et al. 2024).
@@
On the other hand, LLM solvers exhibit strong generalization capabilities when confronted with new COPs and distributions. Recent works leverage LLMs to dynamically generate heuristics or evolve hyper-heuristics (Duflo et al. 2019; Drake et al. 2020), thus facilitating efficient Evolutionary Search (ES) of heuristic spaces without human bias (Yang et al. 2024). Noteworthy endeavors include FunSearch (Romera-Paredes et al. 2024), EoH (Liu et al. 2024a), and ReEvo (Ye et al. 2024). However, existing LLM-based approaches often encounter challenges when handling tasks that require extensive contextual information, typically demonstrating reduced coherence and accuracy in the processing of extended textual descriptions (Yang et al. 2024; Xu et al. 2025). Consequently, these approaches exhibit significant limitations when addressing medium- or large-size instances (e.g., instances involving more than 50 nodes), thus compromising their ability to effectively generalize to real-world VRP instances.
@@ 2 Related Works
Those models learn policies to construct solutions in a step-by-step or one-shot manner: Pointer Network (Ptr-Net) proposed by (Vinyals et al. 2015) employed an encoder-decoder network that generates solutions to VRPs sequentially. Follow-up works utilize reinforcement learning (RL) to improve Ptr-Net’s gradient update to generate better approximate solutions on TSP (Bello et al. 2017) and CVRP (Nazari et al. 2018). More recently, NCO solvers have shifted toward attention-based architectures, starting with the attention model (AM) (Kool et al. 2019), which is versatile on a wide range of COPs. Policy optimization with multiple optima (POMO) (Kwon et al. 2020), has further promoted the capability of the attention-based model on TSP and CVRP. Other constructive frameworks developed from AM and POMO illustrate their broader scalability (Kwon et al. 2021; Choo et al. 2022; Kim et al. 2022a; Chen et al. 2023; Chalumeau et al. 2023; Hottung et al. 2024; Lin et al. 2024).
Recent studies have investigated the severe generalization ability decrement in unseen problem sizes or distributions (Joshi et al. 2006; Liu et al. 2023). In the line of size generalization, there are many attempts to generalize solvers from small instances to larger ones (Lisicki et al. 2020; Kim et al. 2022b; Bdeir et al. 2022; Hou et al. 2023; Son et al. 2023). Unlike approaches focused on scaling, considerable efforts have been devoted to addressing cross-distribution generalization, including works that augment training instances on multi-distribution and multitasks (Goh et al. 2025), and works which synergetically train a backbone model with multi-distribution instances towards a generalizable solver on out-of-distribution tasks (Zhou et al. 2023, 2024).
@@
Existing approaches to improving generalization in combinatorial optimization (CO) take advantage of the linguistic comprehension and domain-specific competence of LLMs via prompt engineering to tackle COPs. These methods can be mainly divided into two mainstreams: 1) LLMs are solution generators. Early explorations in this line of research involve prompting LLMs to solve graph-based COPs (Wang et al. 2023), iteratively refining current heuristics (Iklassov et al. 2024), progressively improving solutions by targeting better objectives (Yang et al. 2024; Liu et al. 2024b; Elhenawy et al. 2024b), and instilling LLMs with visual modalities of problems and solutions (Elhenawy et al. 2024a; Huang et al. 2024). Recently, training LLM for end-to-end CO has been attempted, moving beyond prompting paradigms (Jiang et al. 2025a). 2) LLMs are heuristic designers. Explainable, task-oriented heuristics in the form of codes are iteratively improved throughout automatic heuristic generation and evaluation of Evolution Algorithms (EA), the productivity of which outstrips that of human experts (Romera-Paredes et al. 2024; Liu et al. 2024a; Ye et al. 2024; Tran et al. 2025; Jiang et al. 2025b). Among them, ReEvo (Ye et al. 2024) achieves strong adaptability in many COPs under both white-box and black-box prompt settings. In this paper, we focus mainly on the utilization of LLMs as code designers. Distinct from prior constructive neural solvers and LLM-as-optimizer approaches, our method leverages LLMs to evolve data generators, enabling a novel, generator-based adaptation pipeline that outperforms direct fine-tuning on real data.
@@ 3 Methodology
Representation of Generators. Each generator $h$ is described by three components: 1) a function description defines the format of input and describes valid output, 2) a code implementing generation, and 3) a fitness score. The fitness score, defined as a function $f$ of $h$ (i.e. $f:h\rightarrow\mathbb{R}$), serves as the performance indicator of $h$ on the validation set. An example of an evolved generator that outputs TSP problems is given in Fig.7 in section E of the extended version.
Proxy Evaluation. Following recent work (Ramji et al. 2024), we use an approach that leverages external proxy metrics such as gaps or objective values as the divergence measure $\mathcal{D}$ in Eq.1 to evaluate the quality of LLM-generated outputs, in lieu of direct human or reference-based evaluation. During the LLM-driven evolution, the proxy evaluation is instantiated by measuring the average gap on the validation set, which serves as the proxy metric. The gap is the gap between the objective value found and the best-known optimum. Each generator only fine-tunes the model for a small number of epochs (e.g. 10 epochs for LEHD), and the best average gap on the validation set of each generator during training is taken as the fitness score. Generators with lower fitness scores are retained according to a fixed population size, guiding generator selection without full fine-tuning. This low-fidelity evaluation strategy substantially accelerates the evolution search while presenting reliable indications of the proximity of the generated distributions to the real ones, as is the standard in the neural architecture search and hyperparameter optimization literature.
@@
Evolution Process. In order to balance the effectiveness of evolutionary search and the efficiency of fitness evaluation, we incorporate targeted modifications into ReEvo (Ye et al. 2024) framework. Ours introduces a ranking-based selection mechanism that increases the survival rate of elite generators, thereby accelerating the search process and reducing token costs. In addition, by postponing the selection stage until after mutation, we balance exploration and exploitation, allowing more newly initialized generators to participate in crossover and mutation before selection. Table 5 in Section 5 shows the elevated performance of the evolution framework equipped with these modifications. We perform the workflow in Fig.2 to search for the best generator by repeating the evolution part for N iterations. The whole evolution procedure can be summarized as the following steps:
@@
Phase One. In the first phase of progressive fine-tuning, the best-performing generator of all types evolved collaboratively generate large-scale synthetic data of the same problem size. By training pre-trained models to the disparate node patterns in these diverse, TSPLib-style instances, neural solvers are better equipped to handle challenging patterns in real-world scenarios through preliminary exposure to similar structures. Compared to the proxy evaluation conducted during LLM-driven generator evolution, phase one involves a greater number of training epochs. Additionally, we continuously assess model performance throughout phase one using the 48 validation problems.
@@ 4 Experimental Results
In this section, we evaluate EvoReal on TSPLib95 (Reinelt 1991) and CVRPLib setX (Uchoa et al. 2017).
@@
Baselines. We employ four classic solvers as non-neural baselines: Concorde (Applegate et al. 2006), ORTools, HGS (Vidal 2022) and LKH-3 (Helsgaun 2017). For deep neural solvers, POMO (Kwon et al. 2020), LEHD (Luo et al. 2023), SGBS (Choo et al. 2022), BQ (Drakulic et al. 2023), ELG (Gao et al. 2024), CNF (Zhou et al. 2024) and DIFUSCO (Sun and Yang 2023). Importantly, POMO is trained using reinforcement learning (RL) with policy gradient methods, whereas LEHD employs a supervised learning (SL) scheme that leverages optimal objective values. These models sample solely uniform distributions for training.
@@
Training Setups. Hyperparameters have two sets: one for proxy evaluation and the other for progressive fine-tuning. We choose two representative attention-based neural models for our study, namely POMO and LEHD. For both sets, we leverage the evolved generators to produce TSP or CVRP of size 100 for fine-tuning. In phase two, instances from VRPLib are expanded by a pre-set batch size as a single batch, and they are expanded only once in one epoch. For LEHD, Concorde solver and HGS are adopted to obtain the optimum of TSP and CVRP, respectively. For generator evolution, we use OpenAI model o3 to construct new generators. All evolution pipelines can be conducted on a single NVIDIA RTX 3090 Ti GPU with 24GB memory, whereas the progressive fine-tuning is performed on one NVIDIA RTX A6000 GPU with 48GB memory. The complete hyperparameters are given in section D of the extended version.
Metric and Inference. We compare the results with different metrics with respect to different problem size intervals, including average objective values, average gaps (to the optimal average tour length), and total inference time. In order to monitor the performance of models during the proxy evaluation and progressive fine-tuning, we evaluate the models’ performance at fixed epochs. At each validation point, we calculate the total average gap on the validation set as the performance metric with 8-fold augmentation for POMO or greedy mode for LEHD. The inference time of classic non-neural solvers (CPU) and learning-based models (GPU) are not directly comparable. For POMO, LEHD, SGBS, CNF(3), ELG, DIFUSCO (Ts=50) we report the performance of their pre-trained models with their default settings. For BQ, we reproduced their model with their original setups. For the classic non-neural solvers for TSP, we set a maximum time limit according to problem size $n$ ($T_{max}=1.8*n$ seconds for LKH-3 and ORTools, and $T_{max}=10$ minutes for Concorde). As for non-neural solvers for CVRP, we cite the results from Table 2 of HGS paper (Vidal 2022). For POMO, we provide the results with and without 8 times augmentation (×8 AUG and non aug). With respect to LEHD, the inference results under both greedy mode and RRC (with 50 iterations) are documented.
@@ Appendix A Pre-processing
We follow pre-trained models such as POMO (Kwon et al. 2020) and LEHD (Luo et al. 2023), which randomly sampled uniformly distributed training instances within a unit square. To maintain consistency in data distribution and scale, we also perform the normalization procedure on the node coordinates of TSPLib (Reinelt 1991) and CVRPLib (Uchoa et al. 2017), rescaling them to the unit square before incorporating these datasets into our training pipeline. The Algorithm 1 is applied to the coordinates of all nodes. We also tried other common techniques like min-max normalization, standardization, and without-normalization for fine-tuning, but they either fail to converge or degrade the performance of the pre-trained model. We empirically find that rescaling TSPLib data to match the scale of synthetically generated uniform data, while preserving the original distributional structure, results in a more favorable convergence landscape during fine-tuning with TSPLib instances. For CVRPLib, we also employ the same approach to normalize customer nodes while re-scale the demands with capacities. For LEHD, we normalize the nodes and other constraints of the TSPLib and CVRPLib to the same scale as their training data.
@@
Several prior studies have used statistics to analyze the structural diversity and heterogeneity of TSP instances (Herraiz et al. 2022) or predict optimal tours (Kou et al. 2022). Motivated by these findings, our approach adopts a structure-driven modularization strategy: we categorize TSPLib instances according to statistical features and design specialized generators tailored to each identified category. Specifically, we partition TSPLib using two complementary statistical measures: Fast Fourier Transform (FFT) energy and Nearest-Neighbor ratio (NN-ratio, also known as the Coefficient of Variation, CV). The formal definitions of these two statistics are as follows:
  • FFT Energy: A quantitative statistic that measures the global periodicity or repetitiveness of a spatial point pattern by evaluating the strength of frequency components on its density map. A high FFT energy indicates pronounced global regularity, periodicity, or repeated structures, while a low FFT energy suggests more disordered, chaotic structures. Inspired by classical spectral analysis in signal and spatial statistics (Briggs and Henson 1995; Rajala et al. 2023), FFT energy is widely used to characterize spatial repetition and frequency content in spatial data.
@@
    where $d_{\mathrm{NN}}$ denotes all the nearest neighbor distances in the spatial point set (Clark and Evans 1954; Illian et al. 2008). A low NN-ratio indicates that most points are regularly spaced, as in grid-like or weak-clustered layouts; and a high NN-ratio indicates the presence of clusters or highly varied point spacing. It is a standard, parameter-free measure in spatial statistics to distinguish between regular, random, and clustered point patterns (with self-exclusion, that is, $k=2$ for KNN).
@@
We have also explored other unsupervised machine-learning methods for instance segmentation, such as spectral clustering (Von Luxburg 2007), we found that the classes of most instances (85%) remain the same as those of rule-based segmentation while resulting in trivial performance variation ($5\%$) of the best evolved generators found. This consistency further substantiates the interpretability and robustness of our rule-based classification scheme, indicating that the identified structural discrepancies among different TSPLib distributions are intrinsic rather than heuristic artifacts.